# Deliberation on Priors: Trustworthy Reasoning of Large Language Models on Knowledge Graphs

### v1

@@ 1 Introduction
To equip LLMs with up-to-date, domain-specific knowledge, an emerging line of research [10, 11, 12] has focused on knowledge graph-based retrieval-augmented generation, which aims to enhance response generation through the dynamic retrieval of relevant external knowledge. Current methods [7, 8, 13, 14] employ an end-to-end or step-by-step manner to retrieve and reason on Knowledge Graphs (KGs). The former retrieves top $K$ triples based on the semantic similarity between questions and knowledge facts, while the latter transforms complex questions into multiple subquestions by step-by-step retrieval. After obtaining sufficient knowledge, they utilize LLMs to generate responses directly.
@@
[…]
@@ 2 Preliminary
Table 1: Constraint definitions and samples. Category Definition & Sample Type The question specifies a required type or category for the answer. e.g. What city did Esther live in? Multi- Entity The question demands the answer to satisfy the condition for multiple entities. e.g. Which team owned by Malcolm Glazer has Tim Howard playing for it? Explicit Time The question clearly defines a specific time period or date to be referenced. e.g. Who was the governor of Arizona in 2009 that held his governmental position before 1998? Implicit Time The question implies a temporal frame that the answer should consider. e.g. Who was the Secretary of State when Andrew Jackson was the president? Ordinal The question contains the sorting rule and requires the answer in a specific order. e.g. Lou Seal is the mascot for the team that last won the World Series when?
@@
Path Instantiation. Given a relation path $P$, a instantiation path is called a reasoning path, denoted by $\mathbb{P}=e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}e_{2}\xrightarrow{r_{3}}\cdots\xrightarrow{r_{l}}e_{l}$, where each $e_{i}\in\mathcal{E}$ denotes the $i$-th entity in the path and $r_{i}$ corresponds to the $i$-th relation in the relation path $P$. Here, $\mathcal{E}$ represents the set of all entities in the KG. It should be noted that a relation path $P$ may have multiple instantiations in $\mathcal{G}$.
@@ 3.1 Distillation
Specifically, DP leverages weak supervision signals in the form of question-to-path mappings, which are derived from the annotations within training splits. Given a question $q$ and a topic entity $e_{s}$ identified in $q$, we define a function $\mathcal{M}:\mathcal{Q}\rightarrow 2^{\mathcal{P}}$ that maps $q\in\mathcal{Q}$ to a set of plausible relation paths $\mathcal{P}$, where each $P\in\mathcal{P}$ is a sequence of relations such that traversing from $e_{s}$ via $(r_{1},\dots,r_{l})$ leads to the answer entity $e_{t}$ in the KG $\mathcal{G}=(\mathcal{E},\mathcal{R})$. To extract these paths, our framework first extracts a $k$-hop subgraph $\mathcal{G}_{k}(e_{s})$ centered on $e_{s}$, where $k$ is the maximum reasoning depth allowed by the dataset. Then, the Dijkstra algorithm is employed to identify all shortest relation paths from $e_{s}$ to the ground-truth answer entity $e_{t}$ within $\mathcal{G}_{k}(e_{s})$. These shortest paths constitute the weak supervision set $\mathcal{P}_{w}(q)$ for question $q$, serving as structurally mapping relation exemplars. While not exhaustive, $\mathcal{P}_{w}(q)$ captures plausible multi-hop reasoning trajectories within the KG structure and provides guidance for the LLM to generalize to similar questions. The weak supervision of question-to-path mappings $\mathcal{P}_{w}(q)$ can be defined as follows:
$\mathcal{P}_{w}(q)=\mathcal{M}(q,e_{s},\mathcal{G},k)=\text{ShortestPath}_{\text{Dijkstra}}(\mathcal{G}_{k}(e_{s}),e_{s},e_{t}).$ (1)
@@
To further enhance the reliability of relation path generation, we incorporate a preference optimization stage to explicitly encourage LLMs to prefer semantically coherent and structurally faithful relation paths. Specifically, we construct relation path data $D$ consisting of positive paths and negative paths given question $q$. The negative or undesirable path is synthetically generated from the original weak supervision data $\mathcal{P}_{w}(q)$ through targeted perturbations for the positive path:
@@ H H@1 F1 H H@1 F1 H H@1 F1
Main Result. We compare DP with previous state-of-the-art methods, categorized into three types: SL, ICL, and HL. Based on the results presented in Table 2, we draw the following key insights. First, incorporating deliberation over prior knowledge significantly enhances the reliability of response generation. The results for DP are averaged over three independent runs, with standard deviations reported in Appendix B.3. Our framework consistently achieves new state-of-the-art performance across most datasets, with low variance. Notably, DP surpasses the HL-based method LightPROF by 14.5% in H@1 on the CWQ dataset. Second, existing methods struggle in scenarios where the correct answer must appear in the top-ranked position or where the output is expected to contain accurate answers with high precision. For instance, ToG exhibits a 46.2% gap between its H and F1 scores on the WebQSP dataset. In contrast, DP narrows this gap to approximately 10%, further underscoring its robustness and reliability. Third, we observe that some prior works have misinterpreted evaluation metrics such as H and H@1. According to the code released by [10, 22], certain results reported as H@1 are actually H, potentially leading to unfair comparisons and misleading evaluations. In this work, we rigorously report DP’s performance using H, H@1, and F1 metrics, respectively. Overall, the DP framework demonstrates robust and reliable performance across diverse datasets, validating the effectiveness of incorporating prior knowledge deliberation.
Flexibility Verification. Table 2 also reports the integration experiments, which are designed to evaluate whether DP can enhance the reasoning reliability of various LLMs, including LLaMA3.1-8B, GPT-3.5, GPT-4.0, GPT-4o, and GPT-4.1. The results show that DP consistently improves the reasoning performance of these LLMs across three benchmark datasets. Notably, our framework enables GPT-4o and GPT-4.1 to achieve the best performance on the WebQSP and CWQ datasets, respectively. Overall, these findings highlight the flexibility and effectiveness of DP in enhancing diverse LLMs.
@@ 4.4 Impact of Priors on LLM Reasoning
Table 4: Results of Path Generation (PG) and Constraint Extraction (CE). Setting WebQSP CWQ H F1 H F1 PG (1: 1) 83.0 59.3 81.2 49.8 PG (1: n) 93.0 76.7 94.0 71.1 CE 99.0 90.2 99.0 92.9
@@ 5 Conclusion and Limitation
This work was supported in part by the National Key Research and Development Program of China (2022YFC3303600), the National Natural Science Foundation of China (62306229, 62137002, U22B2019, 62477037, 62293553), the Natural Science Basic Research Program of Shaanxi (2023-JC-YB-593), the Key Research and Development Program of Shaanxi (2024GX-ZDCYL-02-12), the Youth Innovation Team of Shaanxi Universities "Multi-modal Data Mining and Fusion", the Shaanxi Undergraduate and Higher Education Teaching Reform Research Program (23BY195), the Youth Talent Support Program of Shaanxi Science and Technology Association (20240113), and the China Postdoctoral Science Foundation (2024M752585).
@@ B.2.3 Path Selection
- Reasoning Paths: [’Path 1: Baseball -> base.sportbase.sport.played_by_clubs -> Unknown Entity’, Path 2: Deion Sanders -> sports.pro_athlete.teams -> Unknown Entity -> sports.sports_team_roster.team -> Unknown Entity’, Path 3: Deion Sanders -> baseball.baseball_player.batting_stats -> Unknown Entity -> baseball.batting_statistics.team -> Unknown Entity’, Path 4: George Steinbrenner -> sports.sports_team_owner.teams_owned -> Unknown Entity’]

### v2

@@ 1 Introduction
To equip LLMs with up-to-date, domain-specific knowledge, an emerging line of research [10, 11, 12] has focused on knowledge graph-based retrieval-augmented generation, which aims to enhance response generation through the dynamic retrieval of relevant external knowledge. Current methods [7, 8, 13, 14] employ an end-to-end or step-by-step manner to retrieve and reason on Knowledge Graphs (KGs). The former retrieves the top $K$ triples based on the semantic similarity between questions and knowledge facts, while the latter transforms complex questions into multiple subquestions by step-by-step retrieval. After obtaining sufficient knowledge, they utilize LLMs to generate responses directly.
@@
During the online stage, the trained LLM is used to Plan for reasoning and generate faithful candidate relation paths. Subsequently, DP utilizes LLMs to perform relation path selections by evaluating the semantic relevance between the candidate path and the input question. Next, to obtain a complete reasoning path, the framework retrieves topic entities and relations from KGs and instantiates the selected relation path in the Instantiation module, such as the transformation from "{directed, won_award}" to "Ang Lee $\xrightarrow{\texttt{directed}}$ Crouching Tiger $\xrightarrow{\texttt{won\_award}}$ Compound Value Type 2". After obtaining the reasoning path, current frameworks [7, 8, 13, 14, 15, 16] utilize LLMs to generate responses based on them without deliberation. In contrast, to enhance reasoning reliability, DP employs the Introspection module to verify whether the selected reasoning path satisfies the constraints extracted from questions, where the constraints, such as multi-entity and ordinal constraints, are predefined following [17]. Feedback is then provided to the LLM to guide subsequent decisions. If the reasoning path passes the verification, the LLM is instructed to generate a response. Otherwise, the information about which constraints are violated is fed back to the LLM to trigger relation path backtracking.
@@ 2 Preliminary
Table 1: Constraint definitions and samples.
Category Definition & Sample
Type
The question specifies a required type or category for the answer.
e.g. What city did Esther live in?
Multi-
Entity
The question demands the answer to satisfy the condition for multiple entities.
e.g.
Which team owned by Malcolm Glazer has Tim Howard playing for it?
Explicit
Time
The question clearly defines a specific time period or date to be referenced.
e.g.
Who was the governor of Arizona in 2009 that held his governmental position
before 1998?
Implicit
Time
The question implies a temporal frame that the answer should consider.
e.g.
Who was the Secretary of State when Andrew Jackson was the president?
Ordinal
The question contains the sorting rule and requires the answer in a specific order.
e.g.
Lou Seal is the mascot for the team that last won the World Series when?
@@
Path Instantiation. Given a relation path $P$, an instantiation path is called a reasoning path, denoted by $\mathbb{P}=e_{0}\xrightarrow{r_{1}}e_{1}\xrightarrow{r_{2}}e_{2}\xrightarrow{r_{3}}\cdots\xrightarrow{r_{l}}e_{l}$, where each $e_{i}\in\mathcal{E}$ denotes the $i$-th entity in the path and $r_{i}$ corresponds to the $i$-th relation in the relation path $P$. Here, $\mathcal{E}$ represents the set of all entities in the KG. It should be noted that a relation path $P$ may have multiple instantiations in $\mathcal{G}$.
@@ 3.1 Distillation
Specifically, DP leverages weak supervision signals in the form of question-to-path mappings, which are derived from the annotations within training splits. Given a question $q$ and a topic entity $e_{s}$ identified in $q$, we define a function $\mathcal{M}$ that maps $q\in\mathcal{Q}$ to a set of plausible relation paths $\mathcal{P}$, where each $P\in\mathcal{P}$ is a sequence of relations such that traversing from $e_{s}$ via $(r_{1},\dots,r_{l})$ leads to the answer entity $e_{t}$ in the KG $\mathcal{G}=(\mathcal{E},\mathcal{R})$. To extract these paths, our framework first extracts a $k$-hop subgraph $\mathcal{G}_{k}(e_{s})$ centered on $e_{s}$, where $k$ is the maximum reasoning depth allowed by the dataset. Then, the Dijkstra algorithm is employed to identify all shortest relation paths from $e_{s}$ to the ground-truth answer entity $e_{t}$ within $\mathcal{G}_{k}(e_{s})$. These shortest paths constitute the weak supervision set $\boldsymbol{\mathcal{P}}_{w}(q)$} for question $q$, serving as structurally mapping relation exemplars. While not exhaustive, $\boldsymbol{\mathcal{P}}_{w}(q)$} captures plausible multi-hop reasoning trajectories within the KG structure and provides guidance for the LLM to generalize to similar questions. The weak supervision of question-to-path mappings $\boldsymbol{\mathcal{P}}_{w}(q)$} can be defined as follows:
$\boldsymbol{\mathcal{P}}_{w}(q)=\mathcal{M}(q,e_{s},\mathcal{G},k)=\text{ShortestPath}_{\text{Dijkstra}}(\mathcal{G}_{k}(e_{s}),e_{s},e_{t}).$ (1)
@@
To further enhance the reliability of relation path generation, we incorporate a preference optimization stage to explicitly encourage LLMs to prefer semantically coherent and structurally faithful relation paths. Specifically, we construct relation path data $D$ consisting of positive paths and negative paths given question $q$. The negative or undesirable path is synthetically generated from the original weak supervision data $\boldsymbol{\mathcal{P}}_{w}(q)$} through targeted perturbations for the positive path:
@@ H H@1 F1 H H@1 F1 H H@1 F1
[…]
Main Result. We compare DP with previous state-of-the-art methods, categorized into three types: SL, ICL, and HL. Based on the results presented in Table 2, we draw the following key insights. First, incorporating deliberation over prior knowledge significantly enhances the reliability of response generation. The results for DP are averaged over three independent runs. Our framework consistently achieves new state-of-the-art performance across most datasets, with low variance. In contrast, the vanilla LLMs (without DP) exhibit unstable performance and often fail to achieve competitive H@1 or F1 scores, particularly on CWQ and WebQSP, highlighting their limitations in complex reasoning tasks. Notably, DP surpasses the HL-based method LightPROF by 14.5% in H@1 on the CWQ dataset. Second, existing methods struggle in scenarios where the correct answer must appear in the top-ranked position or where the output is expected to contain accurate answers with high precision. For instance, ToG exhibits a 46.2% gap between its H and F1 scores on the WebQSP dataset. In contrast, DP narrows this gap to approximately 10%, further underscoring its robustness and reliability. Third, we observe that some prior works have misinterpreted evaluation metrics such as H and H@1. According to the code released by [10, 22], certain results reported as H@1 are actually H, potentially leading to unfair comparisons and misleading evaluations. In this work, we rigorously report DP’s performance using H, H@1, and F1 metrics, respectively. Overall, the DP framework demonstrates robust and reliable performance across diverse datasets, validating the effectiveness of incorporating prior knowledge deliberation.
Flexibility Verification. Table 2 also reports the integration experiments, which are designed to evaluate whether DP can enhance the reasoning reliability of various LLMs, including LLaMA3.1-8B, GPT-3.5, GPT-4.0, GPT-4o, and GPT-4.1. The results show that DP consistently improves the reasoning performance of these LLMs across three benchmark datasets. Notably, our framework enables GPT-4o and GPT-4.1 to achieve the best performance on the WebQSP and CWQ datasets, respectively. Overall, these findings highlight the flexibility and effectiveness of DP in enhancing diverse LLMs, with lower variance observed across multiple runs, indicating that the improvements are stable and reliable.
@@ 4.4 Impact of Priors on LLM Reasoning
Table 4: Results of Path Generation (PG) and Constraint Extraction (CE).
Setting WebQSP CWQ
H F1 H F1
PG (1: 1) 83.0 59.3 81.2 49.8
PG (1: n) 93.0 76.7 94.0 71.1
CE 99.0 90.2 99.0 92.9
@@ H F1 H F1
4.5 Backtracking Analysis
Table 5: Average number of backtracking steps per question across datasets and models.
Method Dataset Model Backtrack
DP CWQ GPT-3.5 0.10
GPT-4.1 0.42
WebQSP GPT-3.5 0.07
GPT-4.1 0.21
Backtracking refers to the iterative re-execution of relation path selection, instantiation, and introspection when a constraint violation is detected. To explore the practical impact of backtracking, we measure how often it is triggered during test time. Table 5 reports the average number of backtracking steps per question across CWQ and WebQSP. We observe that GPT-4.1 triggers backtracking more frequently than GPT-3.5 (e.g., 0.21 vs. 0.07 on WebQSP), likely due to its stronger instruction-following ability, which enforces stricter constraint checking. This aligns with the performance differences presented in Table 2, suggesting that more reliable backtracking contributes to better overall reasoning quality.
@@ 5 Conclusion and Limitation
This work was supported in part by the National Key Research and Development Program of China (2022YFC3303600), the National Natural Science Foundation of China (62306229, 62137002, U22B2019, 62477037, 62293553), the Natural Science Basic Research Program of Shaanxi (2023-JC-YB-593), the Key Research and Development Program of Shaanxi (2024GX-ZDCYL-02-12), the Youth Innovation Team of Shaanxi Universities "Multi-modal Data Mining and Fusion", the Shaanxi Undergraduate and Higher Education Teaching Reform Research Program (23BY195), the Youth Talent Support Program of Shaanxi Science and Technology Association (20240113), and the China Postdoctoral Science Foundation (2024M752585, 2025T180425).
@@ B.2.3 Path Selection
- Reasoning Paths: [’Path 1: Baseball -> $base.sport$.played_by_clubs -> Unknown Entity’, Path 2: Deion Sanders -> sports.pro_athlete.teams -> Unknown Entity -> sports.sports_team_roster.team -> Unknown Entity’, Path 3: Deion Sanders -> baseball.baseball_player.batting_stats -> Unknown Entity -> baseball.batting_statistics.team -> Unknown Entity’, Path 4: George Steinbrenner -> sports.sports_team_owner.teams_owned -> Unknown Entity’]